OpenStudy (loser66):

A sphere of mass 10kg and radius 0.010m is released from the top of a 1.0meter high inclined plane. What is the speed of the sphere when it reaches the bottom of the inclined plane? the answer is 3.7m/s but I don't know how to get it Please, help

OpenStudy (schrodingers_cat):

Equate energy for the system so you know that Kf + Uf = Ki + Ui 1/2mv^2 +1/2Iw^2 = Mgh 1/2(I/R^2 + M)v^2 = MGh V = (2gh/(1 +(I/MR^2)))^1/2 Hope this Helps :)

OpenStudy (loser66):

cross my fingers, hehehe. I am waiting for help for 1 hour!!!

OpenStudy (loser66):

I got it, Thanks a lot.

OpenStudy (anonymous):

Sorry not to have seen this until today. Glad you got it.

OpenStudy (loser66):

thanks for reply anyway.